OpenStudy (anonymous):

The mass of the first x meters of a thin rod is given by the function m(x) on the indicated interval. Find the linear density function for the rod. Based on what you find, briefly describe the composition of the rod. 1) m(x) = 4x – sin x grams for 0 ≤ x ≤ 6 2) m(x) = 4x grams for 0 ≤ x ≤ 2

OpenStudy (anonymous):

@genius12 can u help?

OpenStudy (anonymous):

@bahrom7893 can u help?

OpenStudy (bahrom7893):

Sorry I have no idea what linear density function is. @Mertsj might now.

OpenStudy (bahrom7893):

I want to say it's just the integral of m(x), but I might be very wrong

OpenStudy (anonymous):

okay

OpenStudy (bahrom7893):

I got it backwards, it's the derivative.

OpenStudy (anonymous):

"Copies and pastes off Wikipedia...." ^^

OpenStudy (mertsj):

yes.

OpenStudy (anonymous):

"without citing..."

OpenStudy (mertsj):

Hunt it up yourselves then.

OpenStudy (anonymous):

Never mind ya'll I think I got it.

OpenStudy (anonymous):

Find dm/dx for each situation. 1) dm/dx = 4 - cos(x) for 0 ≤ x ≤ 6 2) dm/dx = 4 for 0 ≤ x ≤ 2

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

@onegirl

OpenStudy (anonymous):

there is a formula you use hold on let write it up

OpenStudy (anonymous):

\[p(x1) = \lim_{x \rightarrow x1} = f(x) - f(x1)/x - x1 = f'(x1) \]

OpenStudy (anonymous):

wow its so much simpler then it seems to be thank god!